OpenStudy (anonymous):

log(base10) (x-4)+2 x-intercept?

OpenStudy (anonymous):

401/100

OpenStudy (anonymous):

....you probably mean y = logbase 10 (x-4)+2 or f(x).. since otherwise its just an expression and therefore incorrect. Expression != Equation ever. well at x intercept y would be 0. so.. y = log base10 (x-4) + 2 so: -2 = log base 10 (x-4) 10^-2 = 10^ 2(log base 10 (x-4)) 1/100 = x-4 1-100 + 4 = x 401/100 = x